OpenStudy (anonymous):

(Stoichiometry) The annual production of sulfur dioxide from burning coal and fossil fuels, for auto exhaust, and from other activities is about 26 million tons. How much sulfur, present in the original materials, would result in that quantity of SO2? S+O2 --> SO2

OpenStudy (aaronq):

convert 26 million tons of SO2 to moles of SO2 Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.